f(x) = 3x__ and g(x) = __2__
X+2 X+2
Should read 3x over x+2 and 2 Over x+2
Find (f + g) (2)
I know I multiple (x) (2)
Where do I go from here?
What is the domain of f+g?
I’m really unsure about this problem.

Found 2 solutions by Mathtut, solver91311:
Answer by Mathtut(3670)

  

so f+g(x)=(3x)/(x+2)+2/(x+2)=(3x+2)/(x+2)
:
(f+g)(2)=(3(2)+2)/(2+2)=8/4=2
:
the domain is all real numbers except -2 because -2+2 in the denominator is not allowed. you would be dividing by zero which is undefined

Answer by solver91311(24713)






The first thing you need to do is specify which is simply:



Now, to evaluate , just substitute 2 for x in the expression for , thus:



(Note that and and finally )

Since is a rational expression the domain is all real numbers except those numbers that would make the denominator equal zero. In this case, the only number that would make the denominator, is -2.

Hence the domain over the Real numbers, in set builder notation, is